OpenStudy (anonymous):

Solve the equation by graphing: x2 + 5x + 8 = 0. A. {-5, 8} B. {2, 4} C. {-4, 5} D. No solution

OpenStudy (anonymous):

x2 =x^2 right?

OpenStudy (anonymous):

ya thts right

OpenStudy (anonymous):

x^2+2*2.5x+2.5*2.5 = 6.25-8 (x + 2.5)^2 = -1.75 so there is no real (number) solution

OpenStudy (anonymous):

yea i was thinking that too, but just wanted to check to see if i was correct

OpenStudy (anonymous):

thnk u @zzcworld and @helpmenowplease ;)